# Manivel’s semi-group property for Kronecker coefficients, generalized blocks of symmetric groups and Saxl conjecture

### v1

@@ Abstract
Keywords: Saxl conjecture, Generalized $t$-block, Manivel’s semi-group property, Symmetric group.
@@ 1 Introduction
The problem of deciding the positivity of Kronecker coefficients born in quantum information theory [4, 5, 6]. Related to Kronecker positivity problem, there is a fundamental conjecture posed by Heide, Saxl, Tiep and Zalesskii [8]. It states that for every positive integer $n\neq 2,4,9$ there is a complex irreducible character of the symmetric group $\mathfrak{S}_{n}$ whose square contains all irreducible characters of $\mathfrak{S}_{n}$ as constituents. When $n$ is a triangular number, a candidate for this irreducible character of $\mathfrak{S}_{n}$ was suggested by Saxl in 2012: For $k\geq 1$, the staircase partition $\rho_{k}$ of $n:=\binom{k+1}{2}$ is $\rho_{k}=(k,k-1,\dots,1)$. Saxl’s Conjecture says that the Kronecker square $\rho_{k}]^{2}$ contains all irreducible characters of $\mathfrak{S}_{n}$ as constituents. This conjecture has inspired a lot of recent research. For instance, we refer the reader to [2, 9, 15].
Let $n$ be a positive integer. We denote by $\sharp_{n}$ the rectangular partition $(n^{n})\vdash n^{2}$. Suppose that $\alpha=(\alpha_{1},\alpha_{2},\dots,\alpha_{h})\vdash n$. We denote by $\mathrm{Kron}(\alpha)$, the set of all partitions $\beta\vdash n$ satisfying $g(\alpha,\alpha,\beta)\neq 0$. In addition, for every pair $(a,b)$ of non-negative integers, we define $\alpha^{(a,b)}:=(\alpha_{1}+a,\alpha_{2},\dots,\alpha_{h},1^{b})\vdash n+a+b$. To formulate our main results, we require the following definitions.
@@
We now recall Manivel’s semi-group property for Kronecker coefficients [12] which allows us to break partitions down into smaller ones.
@@
(Manivel’s semi-group property [12].) Let $\alpha,\beta,\gamma\vdash n_{1}$, and $\lambda,\mu,\nu\vdash n_{2}$. If both $g(\alpha,\beta,\gamma)>0$ and $g(\lambda,\mu,\nu)>0$ then $g(\lambda+\alpha,\mu+\beta,\nu+\gamma)\geq\mathrm{max}\{g(\lambda,\mu,\nu),g(\alpha,\beta,\gamma)\}$.
@@
Given a positive integer $n$, let $\xi\vdash n$. We start with the following result which is a nice method for detecting some constituents of the Kronecker square $\xi]^{2}$.
Proposition 1.4.
Let $\xi\vdash n$ be a $t$-core partition. If $\alpha\in\mathrm{Kron}(\xi)$, then either
  • a)
    $\alpha$ is a $t$-core partition, or
  • b)
[…]
@@
We let $\mathfrak{S}_{n}$ denote the symmetric group on $n$ letters. We will briefly review some definitions on the representation theory of symmetric groups [10]. A partition $\lambda$ of a positive integer $n$, denoted $\lambda\vdash n$, is defined to be a weakly decreasing sequence $\lambda=(\lambda_{1},\dots,\lambda_{l})$ of positive integers such that the sum $|\lambda|=\sum_{i=1}^{l}\lambda_{i}$ is equal to $n$. The length $l(\lambda)$ of a partition $\lambda\vdash n$ is the number of parts of $\lambda$. For convenience, the notation $(\lambda_{1}^{a_{1}},\dots,\lambda_{r}^{a_{r}})\vdash n$ is used to denote the partition where $\lambda_{i}$’s are the distinct parts that occur with multiplicities $a_{i}$’s. We identify a partition $\lambda$ with its associated Young diagram, that is, the set of cells $\{(i,j)\in\mathbb{N}^{2}_{>0}\,|\,1\leq i\leq l(\lambda),\,j\leq\lambda_{i}\}$. The conjugate or transpose $\lambda^{\prime}$ of $\lambda$ is defined to be equal to the partition obtained from $\lambda$ by reflecting its Young diagram through the $45^{\circ}$ diagonal. Given partitions $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{s})\vdash m$ and $\mu=(\mu_{1},\mu_{2},\dots,\mu_{t})\vdash n$, the partition $\lambda+\mu\vdash m+n$ is defined as follows: Without loss of generality, we can assume that $s\geq t$. The $i$-th part of $\lambda+\mu$ is given by
@@
[10, 2.1.8] Let $\lambda\vdash n$. Then $\lambda^{\prime}]=[(1^{n})][\lambda]$.
@@
(Murnaghan-Nakayama Rule).[16, Theorem 4.10.2] Given positive integers $m$ and $n$, let $\delta\in\mathfrak{S}_{m+n}$ be an $m$-cycle and let $\pi$ be a permutation of the remaining $n$ elements. Then for every $\lambda\vdash m+n$,
@@
Let $\lambda$ be a $t$-core partition and let $\alpha=(\alpha_{1},\alpha_{2},\dots,\alpha_{h})\vdash n$ be $t$-singular. Then for some integer $1\leq i\leq h$, the part $\alpha_{i}$ is divisible by $t$. Using [10, 2.7.40], $\lambda$ is an $\alpha_{i}$-core partition. Hence, $\lambda$ does not have any cell of hook length $\alpha_{i}$. Therefore, it follows from the Murnaghan-Nakayama rule that $\lambda](\alpha)=0$. Conversely, suppose that $\lambda](\alpha)=0$, for every $t$-singular partition $\alpha\vdash n$, but $\lambda$ is not a $t$-core partition. Then $w_{t}(\lambda)\geq 1$. If $\alpha:=(t^{w_{t}(\lambda)},1^{n-t.w_{t}(\lambda)})\vdash n$, then it follows from [10, Corollary 2.7.33] that $\lambda](\alpha)=c[\mathrm{Cor}_{t}(\lambda)]((1^{n-t.w_{t}(\lambda)}))$, for some integer $c\neq 0$. In particular, $\lambda](\alpha)\neq 0$ which is a contradiction as $\alpha$ is $t$-singular. ∎
@@ 3 Proof of our main results
In this section, we wish to prove our main results. We start with a useful method to find constituents of $\rho_{k}]^{2}$, which is inspired by Manivel’s semi-group property.
@@
a) Let $n\geq 2$ be a positive integer. First note that the Kronecker coefficient $g((1^{n}),(1^{n}),(n))=1\neq 0$. It follows from Manivel’s semi-group property that for every $\alpha\in\mathrm{Kron}(\rho_{n-1})$,
@@
b) Let $\lambda,\mu\in\mathrm{Kron}(\rho_{m})$ and $\nu\in\mathrm{Kron}(\sharp_{k-m}).$ Then using Manivel’s semi-group property, we deduce that the Kronecker coefficient
@@
is non-zero. Hence, it again follows by Manivel’s semi-group property that the Kronecker coefficient
@@
Proof of Proposition 1.4. Suppose that $B$ is the $t$-block of the symmetric group $\mathfrak{S}_{n}$ containing the irreducible character $\alpha]$. If $|B|=1$, then as $B$ is a combinatorial $t$-block, we deduce that $\alpha$ is a $t$-core partition. Thus, we can assume that $|B|\geq 2$. Hence, using Lemma 2.4, we get
$\sum_{\pi\in\mathfrak{S}^{(t)}_{n}}[\xi](\pi)[\xi](\pi)[\alpha](\pi)=0$
@@
$g(\xi,\xi,\alpha)=\langle[\xi]^{2},[\alpha]\rangle=\langle[\xi]^{2},[\alpha]\rangle_{\mathfrak{S}^{(t^{\prime})}_{n}}$.
@@
$g(\xi,\xi,\alpha)=\sum_{\lambda\vdash n}g(\xi,\xi,\lambda)\langle[\lambda],[\alpha]\rangle_{\mathfrak{S}^{(t^{\prime})}_{n}}=\sum_{[\lambda]\in B}g(\xi,\xi,\lambda)\langle[\lambda],[\alpha]\rangle_{\mathfrak{S}^{(t^{\prime})}_{n}}$.
Since $|B|\geq 2$, the partition $\alpha$ is not a $t$-core partition and so, it follows from Lemma 2.4 that
@@
$\sum_{[\lambda]\in B\backslash\{[\alpha]\}}g(\xi,\xi,\lambda)\langle[\lambda],[\alpha]\rangle_{\mathfrak{S}^{(t^{\prime})}_{n}}=g(\xi,\xi,\alpha)(1-\langle[\alpha],[\alpha]\rangle_{\mathfrak{S}^{(t^{\prime})}_{n}})$
@@
Note that it follows from [1, Corollary 3.2] that $\sharp_{2^{i}}\in\mathrm{Kron}(\sharp_{2^{i}})$, for every positive integer $s\leq i\leq t$. For the case $m=0$, the result is clear. Thus we can assume that $k\geq 2$. We first claim that $\lambda(t):=\beta_{s}+\sum_{i=s}^{t}(\sharp_{2^{i}}+\alpha_{i})\in\mathrm{Kron}(\rho_{2^{t+1}})$. We prove it by induction on $t$. If $t=s$, then $\lambda(s)=\beta_{s}+\alpha_{s}+\sharp_{2^{s}}$, and so the result follows from Lemma 3.1(b). Now suppose that $t>s$. Given an integer $s\leq r<m-1$, we assume that the claim holds for the case $t=r$ and we try to prove it for $t=r+1$. By induction hypothesis, $\lambda(r)\in\mathrm{Kron}(\rho_{2^{r+1}})$. Hence it follows from Lemma 3.1(b) that $\lambda(r+1)=\lambda(r)+\alpha_{r+1}+\sharp_{2^{r+1}}\in\mathrm{Kron}(\rho_{2^{r+2}})$. Therefore, as $\alpha=\lambda(t)^{(a,b)}$, the result follows from Lemma 3.1(a).
@@
Let $k$ be a positive integer, and let $m:=\lfloor\log_{2}k\rfloor$. Given a positive integer $1\leq i\leq m-1$, suppos that $\alpha_{i}:=((2^{i-1})^{2},1^{2^{2i-1}-2^{i-1}})\vdash\binom{2^{i}+1}{2}$. Then applying [13, Theorem 1.4], $\alpha_{i}$ is of odd degree. Hence, it follows from [2, Theorem 5.2] that $\alpha_{i}\in\mathrm{Kron}(\rho_{2^{i}})$. Therefore, using Theorem 1.5 (a), we have that for every $(a,b)\in G(k,2^{m})$, the Kronecker coefficient $g(\rho_{k},\rho_{k},(\rho_{1}+\sum_{i=0}^{m-1}(\sharp_{2^{i}}+\alpha_{i}))^{(a,b)})$ is non-zero.
@@
Data sharing is not applicable to this article, as no data sets were generated or analysed during the current study.

### v3

# The problem of deciding the positivity of Kronecker coefficients and Saxl conjecture

@@ Abstract
Keywords: Saxl conjecture, Generalized $t$-block, semi-group property, Symmetric group.
@@ 1 Introduction
The problem of deciding the positivity of Kronecker coefficients born in quantum information theory [5, 6, 7]. Related to Kronecker positivity problem, there is a fundamental conjecture posed by Heide, Saxl, Tiep and Zalesskii [9]. It states that for every positive integer $n\neq 2,4,9$ there is a complex irreducible character of the symmetric group $\mathfrak{S}_{n}$ whose square contains all irreducible characters of $\mathfrak{S}_{n}$ as constituents. When $n$ is a triangular number, a candidate for this irreducible character of $\mathfrak{S}_{n}$ was suggested by Saxl in 2012: For $k\geq 1$, the staircase partition $\rho_{k}$ of $n:=\binom{k+1}{2}$ is $\rho_{k}=(k,k-1,\dots,1)$. Saxl’s Conjecture says that the Kronecker square $\rho_{k}]^{2}$ contains all irreducible characters of $\mathfrak{S}_{n}$ as constituents. This conjecture has inspired a lot of recent research. For instance, we refer the reader to [2, 10, 16]. We start with the following result which is a partially answer to the problem of deciding the positivity of Kronecker coefficients.
Proposition 1.1.
Given partitions $\xi,\beta,\alpha\vdash n$, let $g(\xi,\beta,\alpha)\neq 0$. If $\xi$ is a $t$-core partition, then either
  • a)
    $\alpha$ is a $t$-core partition, or
  • b)
    there exists a partition $\nu\neq\alpha$ of $n$ such that $\alpha$ and $\nu$ have the same $t$-core and $g(\xi,\beta,\nu)\neq 0$.
Blasiak [3] gave a positive combinatorial formula for the Kronecker coefficient $g(\lambda,\mu(d),\nu)$ for any partitions $\lambda,\nu$ of $n$ and hook partition $\mu(d):=(n-d,1^{d})$. The following result guaranties the existence of an irreducible constituent indexed by a hook partition $\neq(n),(1^{n})$ in the Kronecker product $\alpha]^{2}$ of a partition $\alpha\neq(n),(1^{n})$ of $n$.
Corollary 1.2.
Let $\alpha\vdash n$. If $\alpha\neq(n),(1^{n})$, then there is a hook partition $\beta\neq(n),(1^{n})$ of $n$ such that $g(\alpha,\alpha,\beta)\neq 0$.
Let $n$ be a positive integer. We denote by $\sharp_{n}$ the rectangular partition $(n^{n})\vdash n^{2}$. Suppose that $\alpha=(\alpha_{1},\alpha_{2},\dots,\alpha_{h})\vdash n$. We denote by $\mathrm{Kron}(\alpha)$, the set of all partitions $\beta\vdash n$ satisfying $g(\alpha,\alpha,\beta)\neq 0$. In addition, for every pair $(a,b)$ of non-negative integers, we define $\alpha^{(a,b)}:=(\alpha_{1}+a,\alpha_{2},\dots,\alpha_{h},1^{b})\vdash n+a+b$. To formulate the next result, we require the following definitions.
@@
We now recall semi-group property for Kronecker coefficients [13] which allows us to break partitions down into smaller ones.
@@
(semi-group property [13].) Let $\alpha,\beta,\gamma\vdash n_{1}$, and $\lambda,\mu,\nu\vdash n_{2}$. If both $g(\alpha,\beta,\gamma)>0$ and $g(\lambda,\mu,\nu)>0$ then $g(\lambda+\alpha,\mu+\beta,\nu+\gamma)\geq\mathrm{max}\{g(\lambda,\mu,\nu),g(\alpha,\beta,\gamma)\}$.
@@
We let $\mathfrak{S}_{n}$ denote the symmetric group on $n$ letters. We will briefly review some definitions on the representation theory of symmetric groups [11]. A partition $\lambda$ of a positive integer $n$, denoted $\lambda\vdash n$, is defined to be a weakly decreasing sequence $\lambda=(\lambda_{1},\dots,\lambda_{l})$ of positive integers such that the sum $|\lambda|=\sum_{i=1}^{l}\lambda_{i}$ is equal to $n$. The length $l(\lambda)$ of a partition $\lambda\vdash n$ is the number of parts of $\lambda$. For convenience, the notation $(\lambda_{1}^{a_{1}},\dots,\lambda_{r}^{a_{r}})\vdash n$ is used to denote the partition where $\lambda_{i}$’s are the distinct parts that occur with multiplicities $a_{i}$’s. We identify a partition $\lambda$ with its associated Young diagram, that is, the set of cells $\{(i,j)\in\mathbb{N}^{2}_{>0}\,|\,1\leq i\leq l(\lambda),\,j\leq\lambda_{i}\}$. The conjugate or transpose $\lambda^{\prime}$ of $\lambda$ is defined to be equal to the partition obtained from $\lambda$ by reflecting its Young diagram through the $45^{\circ}$ diagonal. A Hook partition of $n$ is a partition of the form $\lambda=(n-m,1^{m})$, where $m\leq n$ is a non-negative integer. Given partitions $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{s})\vdash m$ and $\mu=(\mu_{1},\mu_{2},\dots,\mu_{t})\vdash n$, the partition $\lambda+\mu\vdash m+n$ is defined as follows: Without loss of generality, we can assume that $s\geq t$. The $i$-th part of $\lambda+\mu$ is given by
@@
[11, 2.1.8] Let $\lambda\vdash n$. Then $\lambda^{\prime}]=[(1^{n})][\lambda]$.
@@
(Murnaghan-Nakayama Rule).[18, Theorem 4.10.2] Given positive integers $m$ and $n$, let $\delta\in\mathfrak{S}_{m+n}$ be an $m$-cycle and let $\pi$ be a permutation of the remaining $n$ elements. Then for every $\lambda\vdash m+n$,
@@
Let $\lambda$ be a $t$-core partition and let $\alpha=(\alpha_{1},\alpha_{2},\dots,\alpha_{h})\vdash n$ be $t$-singular. Then for some integer $1\leq i\leq h$, the part $\alpha_{i}$ is divisible by $t$. Using [11, 2.7.40], $\lambda$ is an $\alpha_{i}$-core partition. Hence, $\lambda$ does not have any cell of hook length $\alpha_{i}$. Therefore, it follows from the Murnaghan-Nakayama rule that $\lambda](\alpha)=0$. Conversely, suppose that $\lambda](\alpha)=0$, for every $t$-singular partition $\alpha\vdash n$, but $\lambda$ is not a $t$-core partition. Then $w_{t}(\lambda)\geq 1$. If $\alpha:=(t^{w_{t}(\lambda)},1^{n-t.w_{t}(\lambda)})\vdash n$, then it follows from [11, Corollary 2.7.33] that $\lambda](\alpha)=c[\mathrm{Cor}_{t}(\lambda)]((1^{n-t.w_{t}(\lambda)}))$, for some integer $c\neq 0$. In particular, $\lambda](\alpha)\neq 0$ which is a contradiction as $\alpha$ is $t$-singular. ∎
@@ 3 Proof of our main results
In this section, we wish to prove our main results. We start with a useful method to find constituents of $\rho_{k}]^{2}$, which is inspired by semi-group property.
@@
a) Let $n\geq 2$ be a positive integer. First note that the Kronecker coefficient $g((1^{n}),(1^{n}),(n))=1\neq 0$. It follows from semi-group property that for every $\alpha\in\mathrm{Kron}(\rho_{n-1})$,
@@
b) Let $\lambda,\mu\in\mathrm{Kron}(\rho_{m})$ and $\nu\in\mathrm{Kron}(\sharp_{k-m}).$ Then using semi-group property, we deduce that the Kronecker coefficient
@@
is non-zero. Hence, it again follows by semi-group property that the Kronecker coefficient
@@
Proof of Proposition 1.1. Suppose that $B$ is the $t$-block of the symmetric group $\mathfrak{S}_{n}$ containing the irreducible character $\alpha]$. If $|B|=1$, then as $B$ is a combinatorial $t$-block, we deduce that $\alpha$ is a $t$-core partition. Thus, we can assume that $|B|\geq 2$. Using Lemma 2.4, we get
$\sum_{\pi\in\mathfrak{S}^{(t)}_{n}}[\xi](\pi)[\beta](\pi)[\alpha](\pi)=0.$
@@
$g(\xi,\beta,\alpha)=\langle[\xi][\beta],[\alpha]\rangle=\langle[\xi][\beta],[\alpha]\rangle_{\mathfrak{S}^{(t^{\prime})}_{n}}$.
@@
$g(\xi,\beta,\alpha)=\sum_{\lambda\vdash n}g(\xi,\beta,\lambda)\langle[\lambda],[\alpha]\rangle_{\mathfrak{S}^{(t^{\prime})}_{n}}=\sum_{[\lambda]\in B}g(\xi,\beta,\lambda)\langle[\lambda],[\alpha]\rangle_{\mathfrak{S}^{(t^{\prime})}_{n}}$.
Since $|B|\geq 2$ the partition $\alpha$ is not a $t$-core partition and so, it follows from Lemma 2.4 that
@@
$\sum_{[\lambda]\in B\backslash\{[\alpha]\}}g(\xi,\beta,\lambda)\langle[\lambda],[\alpha]\rangle_{\mathfrak{S}^{(t^{\prime})}_{n}}=g(\xi,\beta,\alpha)(1-\langle[\alpha],[\alpha]\rangle_{\mathfrak{S}^{(t^{\prime})}_{n}})$
@@
Proof of Corollary 1.2. If $\alpha$ is a hook partition, then it follows from [17, Theorem 2.1] that $g(\alpha,\alpha,\alpha)$ or $g(\alpha,\alpha,\alpha^{\prime})$ is non-zero. Therefore, without loss of generality, we can assume that $\alpha$ is not a hook partition. At first, let $\alpha\neq\alpha^{\prime}$. It is easily checked that $g(\alpha,\alpha,(n))=1$ and $g(\alpha,\alpha,(1^{n}))=0$. Note that the $n$-block of the symmetric group $\mathfrak{S}_{n}$ is precisely the set of all irreducible characters indexed by hook partitions of $n$. Thus, applying Proposition 1.1, we deduce that there exists a hook partition $\beta\neq(n),(1^{n})$ such that $g(\alpha,\alpha,\beta)\neq 0$. Now, let $\alpha=\alpha^{\prime}=(\alpha_{1},\alpha_{2},\dots,\alpha_{h})$. Set $\alpha^{*}=(\alpha_{2},\alpha_{3},\dots,\alpha_{h})^{\prime}$. Since $\alpha^{*}$ is not self-conjugate, there exists a hook partition $\beta^{*}\neq(n-\alpha_{1}),(1^{n-\alpha_{1}})$ such that $g(\alpha^{*},\alpha^{*},\beta^{*})\neq 0$. Therefore as $g((1^{\alpha_{1}}),(1^{\alpha_{1}}),(\alpha_{1}))=1$, it follows from semi-group property that $g(\alpha,\alpha,\beta^{*}+(\alpha_{1}))\neq 0$. This completes the proof. ∎
@@
Note that it follows from [1, Corollary 3.2] that $\sharp_{2^{i}}\in\mathrm{Kron}(\sharp_{2^{i}})$, for every positive integer $s\leq i\leq t$. For the case $m=0$, the result is clear. Thus, we can assume that $k\geq 2$. We first claim that $\lambda(t):=\beta_{s}+\sum_{i=s}^{t}(\sharp_{2^{i}}+\alpha_{i})\in\mathrm{Kron}(\rho_{2^{t+1}})$. We prove it by induction on $t$. If $t=s$, then $\lambda(s)=\beta_{s}+\alpha_{s}+\sharp_{2^{s}}$, and so the result follows from Lemma 3.1(b). Now suppose that $t>s$. Given an integer $s\leq r<m-1$, we assume that the claim holds for the case $t=r$ and we try to prove it for $t=r+1$. By the induction hypothesis, $\lambda(r)\in\mathrm{Kron}(\rho_{2^{r+1}})$. Hence it follows from Lemma 3.1(b) that $\lambda(r+1)=\lambda(r)+\alpha_{r+1}+\sharp_{2^{r+1}}\in\mathrm{Kron}(\rho_{2^{r+2}})$. Therefore, as $\alpha=\lambda(t)^{(a,b)}$, the result follows from Lemma 3.1(a).
@@
Let $k$ be a positive integer, and let $m:=\lfloor\log_{2}k\rfloor$. Given a positive integer $1\leq i\leq m-1$, suppos that $\alpha_{i}:=((2^{i-1})^{2},1^{2^{2i-1}-2^{i-1}})\vdash\binom{2^{i}+1}{2}$. Then applying [14, Theorem 1.4], $\alpha_{i}$ is of odd degree. Hence, it follows from [2, Theorem 5.2] that $\alpha_{i}\in\mathrm{Kron}(\rho_{2^{i}})$. Therefore, using Theorem 1.6 (a), we have that for every $(a,b)\in G(k,2^{m})$, the Kronecker coefficient $g(\rho_{k},\rho_{k},(\rho_{1}+\sum_{i=0}^{m-1}(\sharp_{2^{i}}+\alpha_{i}))^{(a,b)})$ is non-zero.
@@
Data sharing is not applicable to this article, as no data sets were generated or analyzed during the current study.